OpenStudy (anonymous):

Find the value for sin(x/2) if following conditions hold: tan(x) = -4/3 and x is in quadrant IV a. -sqrt(5)/5 b sqrt(5)/5 c. 2sqrt5)/5 D. -2sqrt(5)/5

OpenStudy (anonymous):

tanx = -4/3 as u know tanx = perpendicular/base so we are going to determine hypotenuse by using pythagoras theorem h^2 = 4^2 + 3^2 h^2 = 16+9 h^2= 25 h=5 is all steps clear?